OpenStudy (anonymous):

If sin 53 degrees is close to 4/5 , which is closest to the length of /NO a. 4cm b. 80cm c. 20cm d. 10cm

OpenStudy (anonymous):

This is what I got 53 degrees and 37 degrees are the approximate measures of the angles of a 3-4-5 right triangle. The legs have lengths 3 and 4, and the hypotenuse is 5. The 4 is opposite the larger angle (53 degrees) and the 3 is opposite the 37 degree angle. sin(53) = opposite/hypotenuse = 4/5 sin(37) = opposite/hypotenuse = 3/5 So I think it is C.

OpenStudy (anonymous):

But I am not sure.

OpenStudy (anonymous):

thank you (:

OpenStudy (anonymous):

Welcome. Was it right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

kk